Find an exact solution to the equation below so that 0 < θ < π/3.
1 = sqrt(3)tan(3θ)
θ = _____________________?

Answer by lwsshak3(11628)

Find an exact solution to the equation below so that 0 < θ < π/3.
1 = sqrt(3)tan(3θ)
**
use x for theta
1=√3tan3x
tan 3x=1/√3
3x=π/6
x=π/18